a couple has 4 children find the probability 3 are girls
Answer by stanbon(75887)

a couple has 4 children find the probability 3 are girls
----
It's the same as the probability of having 1 boy.
---
P(3 girls in 4) = 4C3(1/2)^3(1/2) = 4(1/2)^4 = 4(1/16) = 1/4
=======================
Cheers,
Stan H.